please explain how to solve and simplify this expression. Thank you
0.2 - 0.3(x + 7)+ 0.6 -3

Answer by shrishti(3)

0.2 - 0.3(x + 7)+ 0.6 -3
0.2-0.3x-2.1+0.6-3
-0.3x-4.3
-(0.3x+4.3)=0
- is equal to 0
0.3x+4.3=0
0.3x=-4.3
x=-4.3/0.3